Three tankers contain 403 litres, 434 litres and 465 litres of diesel respectively. Find the maximum capacity of a container that can measure the diesel of the three containers exact number of times.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

HCF of 403, 434, 465 = Maximum capacity of tanker required
403 = 13× 31
434 = 2 × 7 × 31
465 = 3 × 5 × 31
HCF = 31
Hence, a container of 31 litres can measure the diesel of the three containers exact number of times.

Thanks